# NAAQA: A Neural Architecture for Acoustic Question Answering

### v2

@@ Abstract
The goal of the Acoustic Question Answering (AQA) task is to answer a free-form text question about the content of an acoustic scene. It was inspired by the Visual Question Answering (VQA) task. In this paper, based on the previously introduced CLEAR dataset, we propose a new benchmark for AQA, namely CLEAR2, that emphasizes the specific challenges of acoustic inputs. These include handling of variable duration scenes, and scenes built with elementary sounds that differ between training and test set. We also introduce NAAQA, a neural architecture that leverages specific properties of acoustic inputs. The use of 1D convolutions in time and frequency to process 2D spectro-temporal representations of acoustic content shows promising results and enables reductions in model complexity. We show that time coordinate maps augment temporal localization capabilities which enhance performance of the network by $\sim$17 percentage points. On the other hand, frequency coordinate maps have little influence on this task. NAAQA achieves 79.5% of accuracy on the AQA task with $\sim$4 times fewer parameters than the previously explored VQA model. We evaluate the perfomance of NAAQA on an independent data set reconstructed from DAQA. We also test the addition of a MALiMo module in our model on both CLEAR2 and DAQA. We provide a detailed analysis of the results for the different question types. We release the code to produce CLEAR2 as well as NAAQA to foster research in this newly emerging machine learning task.
@@
Acoustic question answering, acoustic reasoning, CNN, auditory scene analysis, CLEAR, Coordconv, Coordinate maps
@@ 1 Introduction
In most studies, reasoning is supported by spatial and symbolic representations in the visual domain [undeft, undefu]. However, reasoning and logic relationships can also be studied via representations of sounds [undefv]. Including the auditory modality in studies on reasoning is of particular interest for research in artificial intelligence, but also has implications in real world applications [undefw]. In [undefx], audio was used in combination with video and depth information to recognize human activities. It was shown that sound can be more discriminative than the corresponding visual cues. As an example, imagine using an espresso machine. Besides possibly a display, all information about the different phases of producing coffee, from grinding the beans, to pressing the powder into the holder and brewing the coffee with high pressure hot water are conveyed by the sounds. Detection of abnormalities in machinery where the moving parts are hidden, or the detection of threatening or hazardous events are other examples of the importance of the audio information for cognitive systems.
@@
The audio modality provides important information that can be leveraged in the context of QA reasoning. Audio allows QA systems to answer relevant questions more accurately, or even to answer questions that are not approachable from the visual domain alone. In [undefy], we introduced the AQA task and proposed a new database (CLEAR) to promote research in AQA. The agent’s goal, in the proposed task, was to answer questions related to acoustic scenes composed by a sequence of elementary musical sounds. The questions foster reasoning on the properties of the elementary sounds and their relative and absolute position in the scene. To build CLEAR, we were inspired by the work of Johnson et al. [undeff] for VQA. Similarly, we tested an architecture built for VQA and based on FiLM layers [undefz] on the newly proposed AQA task. Next, [undefaa] [undefaa] proposed to extend the questions to more acoustically realistic situations by developing a new database called DAQA. To evaluate the results, they proposed the MALiMo network which relies on several FiLM layers.
@@
The works cited above use neural network architectures that are largely inspired by image processing research. However, the structure of acoustic data is fundamentally different from that of visual data. This is illustrated for example in [undefab] where two standard data sets in computer vision (MNIST) and speech technology (Google Speech Commands) are compared via T-SNE [undefac]. A legitimate question is whether it is possible to obtain better results (in terms of accuracy and network complexity) by adapting the first layers of the architectures to take into account intrinsic characteristics of acoustic signals. Even within the AQA domain, the properties of acoustic data may vary significantly depending on the nature of the auditory scenes (e.g. CLEAR vs DAQA). It is, therefore interesting to evaluate the impact of the dataset on system performance.
@@
To answer the above questions, we present a study that evaluates the impact of audio pre-processing, of acoustic feature extraction and of dataset characteristics on the performance neural architectures for AQA. When considering performance, we focus both on accuracy and complexity of the models. The main contributions can be summarized as follows:
@@
  • We introduce CLEAR2 a more challenging version of the CLEAR dataset, which comprises scenes of variable duration and different elementary sounds for the training and test sets,
  • we regenerate a version of the DAQA dataset (DAQA)
  • we evaluate different ways of using a VQA architecture on AQA inputs,
@@
  • we study the effect of time and frequency coordinate maps for acoustic data at different levels in the architecture,
  • we introduce new feature extraction modules motivated by the properties of acoustic data,
  • based on the above points, we propose a highly optimized FiLM-based architecture (NAAQA) that outperforms models that are 5 to 10 times more complex in terms of number of parameters,
  • we test the addition of a MALiMo module (from [undefaa]) into our NAAQA architecture,
  • we compare the performance NAAQA and NAAQA+MALiMo on CLEAR2 and DAQA’ to test generality of the methods,
  • we provide a detailed analysis of our results based on question type to improve interpretability.
@@
The rest of the paper is organized as follows: Section 2 reports on recent related work, Section 3 describes both our CLEAR2 dataset and the DAQA dataset, Section 4 presents the QA models we have tested, Section 5 gives details on the experimental settings, Section 6 presents and discusses the results and, finally, Section 7 concludes the paper. Some extra information is reported in the supplementary material, and referred to in the paper through the corresponding section.
@@ 2 Related Work
This section presents previous research in QA systems and the needs for databases in AQA research.
@@ 2.2 Visual Question Answering (VQA)
Visual Question Answering aims to answer questions based on a visual scene. Several VQA datasets are available to the scientific community [undeff, undefad, undefae, undefaf, undefag, undefah, undefh, undefi, undefai, undefg]. However, designing an unbiased dataset is non-trivial. [undefj] [undefj] observed that the type of questions has a strong impact on the results of neural network based systems which motivated research to reduce the bias in VQA datasets [undefaj, undefak, undefal, undefk, undefl, undeff]. Gathering good labeled data is also non-trivial which induced [undefk] and [undefl] [undefk, undefl] to constrain their work to yes/no questions. To alleviate this problem, [undeff] [undeff] proposed the use of synthetic data for both questions and visual scenes. The resulting CLEVR dataset has been extensively used to evaluate neural networks for VQA applications [undefz, undefam, undefan, undefao, undefap, undefaq] which helped foster research on VQA. To create visual scenes, the authors automated a 3D modelling software. This allows for an unlimited supply of labeled data eliminating the time and effort needed for manual annotations. For the questions, they first manually designed semantic representations for each type of question. These representations describe the reasoning steps needed to answer a question (i.e. “find all cubes $|$ that are red $|$ and metallic”). The semantic representations are then instantiated based on the visual scene composition thus creating a question and an answer for a given scene. This gives complete control over the labelling process.
@@ 2.3 Dababases for AQA
As in VQA, using generated data in the design of AQA datasets has substantial advantages. Data can be automatically annotated which saves time and complexity. The number of training examples that can be generated is only limited by the available computational resources. Controlling the generation process gives a complete understanding of the properties and relations of the objects in a scene. This understanding can be leveraged to reduce bias in the dataset and to generate complex questions and their corresponding answers. The CLEAR dataset  [undefy] has been initially generated using semi-synthetic data. The elementary sounds were recordings of musical notes played by various instruments and players. However, the auditory scenes obtained by combining the elementary sounds had fixed duration. Moreover, the same elementary sounds were used to generate the test and training scenes. The DAQA dataset [undefaa] comprises more complex and less stationary elementary natural sounds coming for example from aircrafts, cars, doors, human speaking, bird singing, dog barking, etc. Although more complex and varied than CLEAR, the evaluation also uses the same elementary sound recordings for training and testing.
@@ 2.4 Convolutional neural network on Audio
Convolutional neural networks (CNN) have dominated the visual domain in recent years. More recently, they have also been applied to a number of problems in the acoustic domains such as acoustic scene classification [undefar, undefas, undefat, undefau], music genre classification [undefav, undefat], instrument classification [undefaw, undefax], sound event classification and localization [undefay] and speech recognition [undefat]. Some authors [undefav, undefaw, undefax, undefau] use intermediate representations such as STFT [undefaz], MFCC [undefaaa] or CQT [undefaab] spectrograms while others work directly with the raw audio signal [undefat, undefar].
Square convolutional and pooling kernels are often used to solve visual task such as VQA, visual scene classification and object recognition [undefaac, undefaad, undefaae, undefaaf]. [undefas, undefaag, undefaah] [undefas, undefaag, undefaah] have successfully used visually motivated CNN with square filters to solve audio related tasks. Time-frequency representations of audio signals are however structured very differently than visual representations. [undefaai] [undefaai] explore the performance of different structures of convolutive kernels when working with music signals classification. They propose the use of 1D convolution kernels to capture time-specific or frequency-specific features. They demonstrate that similar accuracy can be reached using a combination of 1D convolutions instead of 2D convolutions by combining 1D time and 1D frequency filters while using much fewer parameters. They also explore rectangular kernels which capture both time and frequency features at different scales. The impact of such strategies for music classification is still an open question in the context of auditory scene analysis.
Coordinate maps initially proposed in [undefaaj] by [undefaaj] have proven successful for processing visual data in the context of VQA. The method consists in augmenting the visual input with matrices containing numbers in the range -1 to 1 which vary either in the $x$ or in the $y$-dimension. With MALiMo [undefaa], the same strategy is used to indicate the simultaneous relative positions of features in frequency and time. [undefaak] [undefaak] proposed Frequency-Aware convolutions which are equivalent to concatenating coordinate maps only in the frequency axis. The effectiveness of coordinate maps on the time dimension for audio signals have not been studied to the best of our knowledge.
In this study we first evaluate the performance of a network initially designed for the VQA task (Visual FiLM) [undefz] on the AQA task. Then we introduce the NAAQA architecture to leverage specific properties of acoustic inputs. For this architecture, we analyze the influence of separate time and frequency coordinate maps. We then study the impact of adding a MALiMo block into our architecture. Finally, we evaluate our model on the DAQA dataset.
@@ 3.1 CLEAR2
CLEAR2 is an updated version of CLEAR [undefy]. A graphical overview of the generation process is depicted in Figure 1. Each record in the dataset is a unique combination of a scene, a question and an answer.
@@
To build acoustic scenes, we prepared a bank of elementary sounds composed of real musical instrument recordings extracted from the Good-Sounds [undefaal] dataset [^1] Each elementary sound in a scene is characterized by an n-tuple: [Instrument, Brightness, Loudness, Musical Note, Duration, Absolute position in scene, Relative position in scene, Global position]. The Brightness property is computed by using the timbralmodels [undefaam] library. A threshold is used to define the label of the sound (Dark or Bright). The Loudness labels are assigned based on the perceptual loudness as defined by the ITU-R BS.1770-4 international normalization standard [undefaan]. Again, a threshold is used to determine if the sound is Quiet or Loud. All attribute values are listed in Table I as possible answers to the questions explained below.. Differently from CLEAR, in CLEAR2 we make sure that the recordings (players, instruments, microphones) of the elementary sounds are different for the training and test sets. For the training set, the bank comprises 135 unique recordings (compared to 56 in CLEAR) sampled at 48KHz including 6 different instruments (bass, cello, clarinet, flute, trumpet and violin), 12 notes (chromatic scale) and 3 octaves. A different set of 135 recordings of the same instruments recorded using different microphones and players is used to create the test set. The acoustic scenes are built by concatenating between 5 to 15 randomly chosen sounds from the elementary sound bank into a sequence (as opposed to CLEAR which comprised fixed duration scenes). Silence segments of random duration are added in-between elementary sounds. The acoustic scenes are then corrupted by filtering to simulate room reverberation and by adding a white uncorrelated uniform noise. Both the amount of noise and reverberation vary from scene to scene with the goal of reducing overfitting.
@@
The generation process was built with extensibility in mind. Different versions of the dataset with fewer or more objects per scene can be generated by using different parameters for the generation script. It is also possible to modify the elementary sounds bank to generate datasets for AQA in other domains, speech or environmental sounds, for example. The code for generating the dataset is available on GitHub [^2] The link will be made available upon publication.
@@ 3.2 Reproducing DAQA
We were not able to fully recreate the DAQA dataset because it relies on some AudioSet [undefaao] youtube videos that have since been deleted. We were able to retrieve 358 sounds out of the 400 sounds that were used to generate the original dataset. We used these sounds to generate the dataset. Changing the number of elementary sounds also impacts the whole generation process. This dataset is therefore different from the original DAQA and will be referred to as DAQA from now on. Our results are therefore not fully comparable to the ones reported in [undefaa]. A list of all the missing sounds is available in Supplementary G.

### v3

@@ Abstract
The goal of the Acoustic Question Answering (AQA) task is to answer a free-form text question about the content of an acoustic scene. It was inspired by the Visual Question Answering (VQA) task. In this paper, based on the previously introduced CLEAR dataset, we propose a new benchmark for AQA, namely CLEAR2, that emphasizes the specific challenges of acoustic inputs. These include handling of variable duration scenes, and scenes built with elementary sounds that differ between training and test set. We also introduce NAAQA, a neural architecture that leverages specific properties of acoustic inputs. The use of 1D convolutions in time and frequency to process 2D spectro-temporal representations of acoustic content shows promising results and enables reductions in model complexity. We show that time coordinate maps augment temporal localization capabilities which enhance performance of the network by $\sim$17 percentage points. On the other hand, frequency coordinate maps have little influence on this task. NAAQA achieves 79.5% of accuracy on the AQA task with $\sim$4 times fewer parameters than the previously explored VQA model. We evaluate the performance of NAAQA on an independent data set reconstructed from DAQA. We also test the addition of a MALiMo module in our model on both CLEAR2 and DAQA. We provide a detailed analysis of the results for the different question types. We release the code to produce CLEAR2 as well as NAAQA to foster research in this newly emerging machine learning task.
@@
Audio, Question Answering, Reasoning, Temporal Reasoning, CLEAR, Coordconv, Auditory scene analysis
@@ 1 Introduction
In most studies, reasoning is supported by spatial and symbolic representations in the visual domain [undeft, undefu]. However, reasoning and logic relationships can also be studied via representations of sounds [undefv]. Including the auditory modality in studies on reasoning is of particular interest for research in artificial intelligence [undefw], but also has implications in real world applications [undefx]. In [undefy], audio was used in combination with video and depth information to recognize human activities. It was shown that sound can be more discriminative than the corresponding visual cues. As an example, imagine using an espresso machine. Besides possibly a display, all information about the different phases of producing coffee, from grinding the beans, to pressing the powder into the holder and brewing the coffee with high pressure hot water are conveyed by the sounds. Detection of abnormalities in machinery where the moving parts are hidden, or the detection of threatening or hazardous events are other examples of the importance of the audio information for cognitive systems.
@@
The audio modality provides important information that can be leveraged in the context of QA reasoning. Audio allows QA systems to answer relevant questions more accurately, or even to answer questions that are not approachable from the visual domain alone. In [undefz], we introduced the AQA task and proposed a new database (CLEAR) to promote research in AQA. The agent’s goal, in the proposed task, was to answer questions related to acoustic scenes composed by a sequence of elementary musical sounds. The questions foster reasoning on the properties of the elementary sounds and their relative and absolute position in the scene. To build CLEAR, we were inspired by the work of Johnson et al. [undeff] for VQA. Similarly, we tested an architecture built for VQA and based on FiLM layers [undefaa] on the newly proposed AQA task. [undefab] [undefab] later proposed to extend the questions to more acoustically realistic situations by developing a new database called DAQA. To evaluate the results, they proposed the MALiMo network which relies on several FiLM layers.
@@
The works cited above use neural network architectures that are largely inspired by image processing research. However, the structure of acoustic data is fundamentally different from that of visual data. This is illustrated for example in [undefac] where two standard data sets in computer vision (MNIST) and speech technology (Google Speech Commands) are compared via T-SNE [undefad]. A legitimate question is whether it is possible to obtain better results (in terms of accuracy and network complexity) by adapting the first layers of the architectures to take into account intrinsic characteristics of acoustic signals. Even within the AQA domain, the properties of acoustic data may vary significantly depending on the nature of the auditory scenes (e.g. CLEAR vs DAQA). It is, therefore interesting to evaluate the impact of the dataset on system performance.
@@
To answer the above questions, we present a study that evaluates the impact of audio pre-processing, of acoustic feature extraction and of dataset characteristics on the performance neural architectures for AQA. When considering performance, we focus both on accuracy and complexity of the models. We provide a detailed analysis of our results based on question type to improve interpretability. The main contributions can be summarized as follows:
@@
  • We introduce CLEAR2 a more challenging version of the CLEAR dataset, which comprises scenes of variable duration and different elementary sounds for the training and test sets.
  • We propose a highly optimized FiLM-based architecture (NAAQA) inspired by VQA tasks containing new feature extraction modules that are tailored to acoustic inputs.
@@
  • We study the effect of time and frequency coordinate maps for acoustic data at different levels in the architecture.
  • We evaluate the generality of the methods by testing NAAQA on a regenerated a version of the DAQA dataset (DAQA) and by adding a MALiMo module (from [undefab]) into our NAAQA architecture.
  • We provide a detailed analysis of our experimental results that helps interpretability of the model.
On the CLEAR2 dataset NAAQA outperforms the VQA baseline (which is 4 times more complex in terms of number of parameters) by 17.2 percent points in the accuracy score.
@@
The rest of the paper is organized as follows: Section 2 reports on recent related work, Section 3 describes both our CLEAR2 dataset and the DAQA dataset, Section 4 presents the QA models we have tested, Section 5 gives details on the experimental settings, Section 6 presents and discusses the results and, finally, Section 7 concludes the paper. Some extra information is reported in the supplementary material.
@@ 2 Related Work
This section presents previous research in QA systems including data generation and modeling.
@@ 2.2 Visual Question Answering (VQA)
Visual Question Answering aims to answer questions based on a visual scene. Several VQA datasets are available to the scientific community [undeff, undefae, undefaf, undefag, undefah, undefai, undefh, undefi, undefaj, undefg]. However, designing an unbiased dataset is non-trivial. [undefj] [undefj] observed that the type of questions has a strong impact on the results of neural network based systems which motivated research to reduce the bias in VQA datasets [undefak, undefal, undefam, undefk, undefl, undeff]. Gathering good labeled data is also non-trivial which induced [undefk] and [undefl] [undefk, undefl] to constrain their work to yes/no questions. To alleviate this problem, [undeff] [undeff] proposed the use of synthetic data for both questions and visual scenes. The resulting CLEVR dataset has been extensively used to evaluate neural networks for VQA applications [undefaa, undefan, undefao, undefap, undefaq, undefar] which helped foster research on VQA. To create visual scenes, the authors automated a 3D modelling software. This allows for an unlimited supply of labeled data eliminating the time and effort needed for manual annotations. For the questions, they first manually designed semantic representations for each type of question. These representations describe the reasoning steps needed to answer a question (i.e. “find all cubes $|$ that are red $|$ and metallic”). The semantic representations are then instantiated based on the visual scene composition thus creating a question and an answer for a given scene. This gives complete control over the labelling process.
@@ 2.3 Dababases for AQA
As in VQA, using generated data in the design of AQA datasets has substantial advantages. Data can be automatically annotated which saves time and complexity. The number of training examples that can be generated is only limited by the available computational resources. Controlling the generation process gives a complete understanding of the properties and relations of the objects in a scene. This understanding can be leveraged to reduce bias in the dataset and to generate complex questions and their corresponding answers. The CLEAR dataset  [undefz] has been initially generated using semi-synthetic data. The elementary sounds were real recordings of musical notes played by various instruments and players. The auditory scenes were obtained by concatenating these elementary sounds in different combinations. The data set had two main limitations: scenes had fixed duration, and the same elementary sounds were used to generate the test and training scenes (although test and training scenes were different). The DAQA dataset [undefab] comprises more complex and less stationary elementary natural sounds coming for example from aircrafts, cars, doors, human speaking, bird singing, dog barking, etc. Although more complex and varied than CLEAR, the evaluation also uses the same elementary sound recordings for training and testing.
@@ 2.4 Convolutional neural network on Audio
Convolutional neural networks (CNN) have dominated the visual domain in recent years. More recently, they have also been applied to a number of problems in the acoustic domains such as acoustic scene classification [undefas, undefat, undefau, undefav], music genre classification [undefaw, undefau], instrument classification [undefax, undefay], sound event classification and localization [undefaz] and speech recognition [undefau]. Some authors [undefaw, undefax, undefay, undefav] use intermediate representations such as STFT [undefaaa], MFCC [undefaab] or CQT [undefaac] spectrograms while others work directly with the raw audio signal [undefau, undefas].
Square convolutional and pooling kernels are often used to solve visual task such as VQA, visual scene classification and object recognition [undefaad, undefaae, undefaaf, undefaag]. [undefat, undefaah, undefaai] [undefat, undefaah, undefaai] have successfully used visually motivated CNN with square filters to solve audio related tasks. Time-frequency representations of audio signals are however structured very differently than visual representations. [undefaaj] [undefaaj] explore the performance of different structures of convolutive kernels when working with music signals classification. They propose the use of 1D convolution kernels to capture time-specific or frequency-specific features. They demonstrate that similar accuracy can be reached using a combination of 1D convolutions instead of 2D convolutions by combining 1D time and 1D frequency filters while using much fewer parameters. They also explore rectangular kernels which capture both time and frequency features at different scales. The impact of such strategies for music classification is still an open question in the context of auditory scene analysis.
Coordinate maps initially proposed in [undefaak] by [undefaak] have proven successful for processing visual data in the context of VQA. The method consists in augmenting the visual input with matrices containing numbers in the range -1 to 1 which vary either in the $x$ or in the $y$-dimension. With MALiMo [undefab], the same strategy is used to indicate the simultaneous relative positions of features in frequency and time. [undefaal] [undefaal] proposed Frequency-Aware convolutions which are equivalent to concatenating coordinate maps only in the frequency axis. The effectiveness of coordinate maps on the time dimension for audio signals have not been studied to the best of our knowledge.
In this study we first evaluate the performance of a network initially designed for the VQA task (Visual FiLM) [undefaa] on the AQA task, using the CLEAR2 data set. Then we introduce the NAAQA architecture to leverage specific properties of acoustic inputs. For this architecture, we analyze the influence of separate time and frequency coordinate maps. We then study the impact of adding a MALiMo block into our architecture. Finally, we evaluate our model on the DAQA dataset.
@@ 3.1 CLEAR2
CLEAR2 is an updated version of CLEAR [undefz]. A graphical overview of the generation process is depicted in Figure 1. Each record in the dataset is a unique combination of a scene, a question and an answer.
@@
To build acoustic scenes, we prepared a bank of elementary sounds composed of real musical instrument recordings extracted from the Good-Sounds [undefaam] dataset [^1] Each elementary sound in a scene is characterized by an n-tuple: [Instrument, Brightness, Loudness, Musical Note, Duration, Absolute position in scene, Relative position in scene, Global position]. The Brightness property is computed by using the timbralmodels [undefaan] library. A threshold is used to define the label of the sound (Dark or Bright). The Loudness labels are assigned based on the perceptual loudness as defined by the ITU-R BS.1770-4 international normalization standard [undefaao]. Again, a threshold is used to determine if the sound is Quiet or Loud. All attribute values are listed in Table I as possible answers to the questions explained below.. Differently from CLEAR, in CLEAR2 we make sure that the recordings (players, instruments, microphones) of the elementary sounds are different for the training and test sets. For the training set, the bank comprises 135 unique recordings (compared to 56 in CLEAR) sampled at 48KHz including 6 different instruments (bass, cello, clarinet, flute, trumpet and violin), 12 notes (chromatic scale) and 3 octaves. A different set of 135 recordings of the same instruments recorded using different microphones and players is used to create the test set. The acoustic scenes are built by concatenating between 5 to 15 randomly chosen sounds from the elementary sound bank into a sequence (as opposed to CLEAR which comprised fixed duration scenes). Silence segments of random duration are added in-between elementary sounds. The acoustic scenes are then corrupted by filtering to simulate room reverberation and by adding a white uncorrelated uniform noise. Both the amount of noise and reverberation vary from scene to scene with the goal of increasing variability in the data.
@@
The generation process was built with extensibility in mind. Different versions of the dataset with fewer or more objects per scene can be generated by using different parameters for the generation script. It is also possible to modify the elementary sounds bank to generate datasets for AQA in other domains, speech or environmental sounds, for example. The code for generating the dataset is available on GitHub [^2] https://github.com/NECOTIS/CLEAR-AQA-Dataset-Generator. Pre-generated version of the dataset is available both on IEEE Dataport [^3] https://dx.doi.org/10.21227/7x26-a025 and HuggingFace [^4] https://huggingface.co/datasets/J3romee/CLEAR.
@@ 3.2 Reproducing DAQA
We were not able to fully recreate the DAQA dataset because it relies on some AudioSet [undefaap] YouTube videos that have since been deleted. We were able to retrieve 358 sounds out of the 400 sounds that were used to generate the original dataset. We used these sounds to generate the dataset. Changing the number of elementary sounds also impacts the whole generation process. This dataset is therefore different from the original DAQA and will be referred to as DAQA from now on. Our results are therefore not fully comparable to the ones reported in [undefab]. A list of all the missing sounds is available in Supplementary Materials.